# ISWC’24: International Semantic Web Conference, 2024, Baltimore, USA

### v3

@@ 1 Motivation
The conference track is one of the Ontology Alignment Evaluation Initiative (OAEI) [7] campaigns that evaluates OM tasks in conference organisation. It includes several reference alignments between different ontologies related to research conference management. For example, “http://cmt#Chairman” and “http://conference#Chair” form a pair of valid match in the CMT-Conference alignment. LLMs can experience hallucinations when they mistakenly categorise a pair of matched entities to be unmatched entities. This is oftentimes because it is very difficult for LLMs to understand polysemous words, such as where the word “chair” refers to the person who chairs the conference, and not to the furniture used at the conference. LLM hallucinations can also occur when one of the matched entities is mismatched with other entities. For example, LLMs may consider “conference chair” to have a similar meaning to “chair” in the context of a research conference. Therefore, “http://conference#Chair” can be mismatched to “http://cmt#ConferenceChair”. Both the former case of missing true mappings and the latter case of generating false mappings will have a negative impact on the matching performance. It seems feasible to mitigate LLM hallucinations using fine-tuning, but there are few datasets that can be used to provide baseline references.
@@ 2.1 Dataset Construction
For each mapping $(e1,e2)\in R_{oaei}$ we consider that it is not an LLM hallucination if there exists $(e1^{\prime},e2^{\prime})\in A_{llm}$ such that $e1=e1^{\prime}$ and $e2=e2^{\prime}$. We categorise LLM mapping errors in three ways (see Table 1 for details):
@@
In practice, the notions of relevant and intended are implemented by the chosen matching assessment process. We use an LLM-based evaluator for matching assessment, and directly query the LLM for relevance and superclass and subclass relationships. We are considering updating our LLM-based evaluator to a multi-agent negotiation and human-in-the-loop validation to enhance its precision. The evaluation produced by the matching assessment process will be used to generate the OAEI-LLM benchmark that can be used for comparison with alternative LLM-based matchers.
@@ 2.2 Schema Extension
We extend the current EDOAL mapping schema to record the new information related to LLM hallucinations. This extension may not be required if one could use a more flexible mapping format like the Simple Standard for Sharing Ontological Mappings (SSSOM) [9] for OAEI tracks. For each mapping, different types of hallucinations can occur in different LLMs. Each of them is recorded in the tag <hallucination>. In the extended version of the example (highlighted in blue), the first segment depicts the LLM hallucinations that occurred in LLM-X, which is a case of a mapping missing from the LLM. The second segment illustrates the LLM hallucinations that occurred in LLM-Y, indicating an incorrect mapping to the entity “conference chair”. The matching assessment classifies this mapping as a “disputed mapping”.
@@ 3.1 Benchmarking LLMs for OM Tasks
3.2 A Dataset for Fine-tuning LLMs Used in OM Tasks
@@ 4 Limitations
In our work to date we use a moderately simple LLM-based evaluator to perform the matching assessment. The assessment may vary depending on the LLM used. A more sophisticated matching assessment could use a voting process by multiple LLMs or human expert judgement to generate more reliable OAEI-LLM benchmark datasets.
@@ Acknowledgements.
The authors thank the reviewers for providing valuable feedback to improve the manuscript. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v4

@@ 1 Motivation
The conference track is one of the Ontology Alignment Evaluation Initiative (OAEI) [7] campaigns that evaluates OM tasks in conference organisation. It includes several reference alignments between different ontologies related to research conference management. For example, “http://cmt#Chairman” and “http://conference#Chair” form a pair of valid match in the CMT-Conference alignment. LLMs can experience hallucinations when they mistakenly categorise a pair of matched entities to be unmatched entities. This often happens because it is very difficult for LLMs to understand polysemous words, such as where the word “chair” refers to the person who chairs the conference, and not to the furniture used at the conference. LLM hallucinations can also occur when one of the matched entities is mismatched with other entities. For example, LLMs may consider “chair” in the context of a research conference to have a similar meaning to “conference chair”. Therefore, “http://conference#Chair” can be mismatched to “http://cmt#ConferenceChair”. Both the former case of missing true mappings and the latter case of generating false mappings will have a negative impact on the matching performance. It seems feasible to mitigate LLM hallucinations using fine-tuning, but there are few datasets that can be used to provide baseline references.
@@ 2.1 Dataset Construction
For each mapping $(e1,e2)\in R_{oaei}$, we consider that it is not an LLM hallucination if there exists $(e1^{\prime},e2^{\prime})\in A_{llm}$ such that $e1=e1^{\prime}$ and $e2=e2^{\prime}$. We categorise LLM mapping errors in three ways (Table 1):
@@
In practice, the notions of relevant and intended are implemented by the chosen matching assessment process. We use an LLM-based evaluator for matching assessment, and directly query the LLM for relevance and superclass and subclass relationships. At this time we do not recommend a particular LLM to be used. We are considering updating our LLM-based evaluator to a multi-agent negotiation and human-in-the-loop validation to enhance its precision. The evaluation produced by the matching assessment process will be used to generate the OAEI-LLM benchmark that can be used for comparison with alternative LLM-based matchers.
@@ 2.2 Schema Extension
We extend the current EDOAL mapping schema to record the new information related to LLM hallucinations. This extension may not be required if one could use a more flexible mapping format like the Simple Standard for Sharing Ontological Mappings (SSSOM) [9] for OAEI tracks. For each mapping, different types of hallucinations can occur in different LLMs. Each of them is recorded in the tag <hallucination>. In the extended version of the example (highlighted in blue), the first segment depicts the LLM hallucinations that occurred in gpt-4o-mini [10], which is a case of a mapping missing from the LLM. The second segment illustrates the LLM hallucinations that occurred in llama-3-8b [11], indicating an incorrect mapping to the entity “conference chair”. The matching assessment classifies this mapping as a “disputed mapping”.
@@ 4 Limitations
In our work to date we use a moderately simple LLM-based evaluator to perform the matching assessment. The assessment may vary depending on the LLM used. A more sophisticated matching assessment could use a voting process by multiple LLMs or human expert judgment to generate more reliable OAEI-LLM benchmark datasets.
@@ 5 Further Work
Supplemental Material Statement
The slide presentation can be found in the uploaded file package.
@@ Acknowledgements.
The authors thank the reviewers for providing valuable feedback during the review process. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.